# Agentless [Uncaptioned image] : Demystifying LLM-based Software Engineering Agents

### v1

@@ 4 Evaluation
We now look at the impact of our different repair setups on the final performance. Table 3 shows the different ways we can generate or select the final patch for submission. Starting with just generating a single sample (i.e., using greedy decoding), Agentless can achieve 70 correct fixes while costing an average of \$0.11 dollars per bug (total cost, including localization). We note that even with this simple patch generation step, Agentless can already beat the majority of the prior open-source agent-based approaches (with more than 4X in cost reduction). We can further improve performance to 78 fixes by sampling the LLM multiple times and selecting a patch using majority voting. Finally, the full Agentless performance is achieved by further applying filtering to select only the patches which can successfully pass the existing regressions tests. Since we sample multiple patches per each issue, we also observe that the total number of possible issues that Agentless can solve when using all samples is 123 (41.0%). This shows a high upper bound for the repair potential of Agentless with future work being better re-ranking and selection techniques to further improve the overall performance.
@@ 5.1 Problem Classification
1) Description quality. We first inspect whether each issue description contains sufficient information to perform the desired task. Figure 5(a) shows the distribution of each category: (i) contains enough information in natural language, (ii) contains reproducible failure example, (iii) contains partially reproducible example, and (iv) does not contain enough information.We observe that while a majority of the tasks in SWE-bench Lite contains sufficient information, with many having some small failure examples to showcase the bug, there is a non-trivial percentage (9.3%) of problems which do not contain enough information. Such problems include those that require implementing a new function with a specific name or adding an error message with a specific string that was not provided in the problem description. [^2] These types of problems still exist in the benchmark despite claims that they have been completely removed by the filtering process according to SWE-bench Lite. This means the test will fail if the function name or error message string does not match exactly, even if the underlying functionality is correctly implemented. Another example of insufficient information are problems that have multiple different interpretations on how to solve the issue, and only a subset of them can pass the ground truth test. For instance, the issue description will outline two possible solutions suggestions with only one of them aligned well with developer intention. Implementing the other proposed solution suggestion will lead to test failure. This highlights the necessity to further sanitize/improve SWE-bench Lite where these problems with uninformative descriptions shall be further excluded.
@@
2) Solution in description. We also examine whether the solution or steps to solve the problem are already provided in the issue description. Figure 5(b) shows the breakdown of our categories: (i) no solution or steps provided, (ii) partial solution provided (e.g., some steps in natural language), (iii) complete solution provided (e.g., complete steps in natural language), (iv) exact patch provided, and (v) misleading solution or steps. Interestingly, we observe that 4.3% of issues contain the exact ground truth patch in the issue description, while an additional 10.0% of issues describe the exact steps required to come up with the correct solution. This shows that certain problems in SWE-bench Lite can be much easier to solve since they provide the solution either in exact code snippets or natural language. Furthermore, we also observe 4.3% of issues contain proposed solution or steps in the issue description that do not reflect the ground truth patch introduced by the developers. This further highlights potential issues with the benchmark, as these discrepancies can mislead tools to generate incorrect solutions simply by following the issue description.
@@
3) Location information. We further check if the issues description contains the correct location information. We divide the granularity into line, function, and file level locations. Our categories are: (i) exact locations in natural language, (ii) exact locations provided in failure stack traces, iii) related keywords in the issue description that can be used to search for the location, and (iv) not provided. We first observe that only in very few cases (<<10%), the issue provides the exact lines needed to fix the bug. However, this number increases as we increase the granularity to functions and files where we found that around half of the issues already provide the location of the file needed to be edited in the description. To repair a bug or introduce a new feature, finding the location to make the edit is extremely important. As such, we leverage this classification and focus our later analysis on the effect the provided location has on the repair performance of Agentless and baseline approaches.
@@ 5.2 SWE-bench Lite-$S$
Alibaba Lingma Agent lin [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o+ [Uncaptioned image] Claude-3.5 99 (33.00%) 1 87 (34.52%) 1
Factory Code Droid fac [2024] [Uncaptioned image] NA 94 (31.33%) 2 82 (32.54%) 2
AutoCodeRover-v2 aut [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o 92 (30.67%) 3 79 (31.35%) 3
CodeR Chen et al. [2024] [Uncaptioned image] [Uncaptioned image] GPT-4 85 (28.33%) 4 72 (28.57%) 4
IBM Research Agent-101 ibm [2024] [Uncaptioned image] NA 80 (26.67%) 6 66 (26.19%) 7
OpenCSG StarShip ope [2024a] [Uncaptioned image] [Uncaptioned image] GPT-4 71 (23.67%) 9 57 (22.62%) 9
Bytedance MarsCode mar [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o 76 (25.33%) 8 63 (25.00%) 8
Amazon Q Developer ama [2024] [Uncaptioned image] NA 61 (20.33%) 11 52 (20.63%) 10*
RepoUnderstander Ma et al. [2024] [Uncaptioned image] [Uncaptioned image] GPT-4 64 (21.33%) 10 52 (20.63%) 10*
Aider Gauthier [2024] [Uncaptioned image] GPT-4o+ [Uncaptioned image] Claude-3 79 (26.33%) 7 67 (26.59%) 6
AutoCodeRover Zhang et al. [2024] [Uncaptioned image] GPT-4 57 (19.00%) 12 46 (18.25%) 12
Claude-3
35 (11.67%) 16 27 (10.71%) 16
SWE-agent Yang et al. [2024] [Uncaptioned image] GPT-4 54 (18.00%) 13 42 (16.67%) 14
[Uncaptioned image] GPT-4 50 (16.67%) 15 41 (16.27%) 15
OpenDevin ope [2024b] [Uncaptioned image] GPT-4o 52 (17.33%) 14 45 (17.86%) 13
Claude-3
13 (4.33%) 17 10 (3.97%) 17
[Uncaptioned image] GPT-4 8 (2.67%) 19 5 (1.98%) 19
@@
9 (3.00%) 18 6 (2.38%) 18
RAG Jimenez et al. [2024a] [Uncaptioned image] ChatGPT 1 (0.33%) 20 0 (0.00%) 20
Agentless [Uncaptioned image] [Uncaptioned image] GPT-4o 82 (27.33%) 5 71 (28.17%) 5
Building on the above problem classifications, in the following evaluation section, we will more rigorously compare and contrast Agentless and existing work. Specifically, we focus on a subset of the problems in SWE-bench Lite after removing the problems that contain the exact patch in the problem description, misleading solutions, or do not provide enough information in the original issue description. This eliminates the less reasonable problems and normalizes the difficulty level of the benchmark. For future work, we hope to work with the maintainers and contribute to the SWE-bench Lite benchmark by fixing these unreasonable problems to add additional information as well as removing exact ground truth patches in the problem descriptions. However, as we are not able to run commercial tools ourselves on the modified problems, we simply exclude the problematic problems in the below evaluation. We refer to our subset of 252 problems as SWE-bench Lite-$S$.
@@
Using the classification results, we further examine the types of problems that are solved by Agentless and prior approaches on SWE-bench Lite-$S$. Figure 6 shows the solve rate of various top-performing open-source and closed-source approaches across the different categories of problems. We first examine if having code examples to reproduce the error in the issue description can help the LLM better solve the issue in Figure 6(a). Surprisingly, we found that the solve rate of all prior approaches drop when evaluated on the problems with reproducible code examples. Many agent-based approaches Yang et al. [2024]; ope [2024b]; Chen et al. [2024] attempt to first reproduce the error, however, this may not improve performance even on problems with already provided reproducible examples. This shows that there are still room for further improvement specifically on reproducing error-triggering tests. Next, we look at the effect of ground truth patch/solution in the issue description. Figure 6(b) shows the expected result where all selected techniques perform better on issues that already provide solution steps in natural language. Furthermore, in Figure 6(c), we examine the solve rate with respect to the location information provided in the issues description. Unsurprisingly, we found that the highest solve rates are on problems where the location is provided in natural language followed by stack traces. The most difficult problems are those that do not contain any clues about the location of the issue in the description. We observe that compared with closed-source approaches, Agentless performs comparably when the location is provided in either natural language, stack trace, or keywords. However, the closed-source agent tools perform better compared to Agentless in the case where no location clue is provided. This highlights an advantage of agent-based tools in solving these more complex problems where they are able to use complex code search tools. This represents potential future work for Agentless to target and further improve these types of problems.
6 Related Work
LLMs for code. LLMs have become the default choice for various coding tasks, due to the impressive results achieved by LLMs in both code generation and understanding Chen et al. [2021]. Developers and researchers have applied on software engineering tasks, such as program synthesis Patton et al. [2024]; Chen et al. [2021]; Li et al. [2022a]; Iyer et al. [2018], code translation Pan et al. [2024]; Roziere et al. [2020, 2021], program repair Xia et al. [2023]; Xia & Zhang [2023]; Monperrus [2018]; Kolak et al. [2022]; Bouzenia et al. [2024], and test generation Deng et al. [2023]; Xia et al. [2024]; Deng et al. [2024]; Lemieux et al. [2023]; Kang et al. [2023]. Apart from using general-purpose LLMs, code-specific LLMs have been built by further training LLMs using large amounts of open-source code snippets. Examples of code LLMs include Codex Chen et al. [2021], CodeLlama Rozière et al. [2023], StarCoder Li et al. [2023]; Lozhkov et al. [2024], DeepSeek-Coder Guo et al. [2024], etc. Furthermore, researchers have also developed instruction-following code-specific LLMs using instruction-tuning methods. Examples of such LLMs include CodeLlama-Inst Rozière et al. [2023], DeepSeek-Coder-Inst Guo et al. [2024], WizardCoder Luo et al. [2023], Magicoder Wei et al. [2023], and OpenCodeInterpreter Zheng et al. [2024].
Benchmarking for LLM-based coding tasks. To evaluate the capability of LLMs on code, various benchmark has been proposed. HumanEval Chen et al. [2021] and MBPP Austin et al. [2021] are two of the most widely-used handcrafted code generation benchmarks complete with test cases to check for the correctness of LLM outputs. Furthermore, other benchmarks have been proposed with more robust test Liu et al. [2023a], additional programming languages Zheng et al. [2023]; Cassano et al. [2023], and other programming domains Jain et al. [2024]; Hendrycks et al. [2021]; Li et al. [2022b]; Lai et al. [2023]; Yin et al. [2022].
More recently, instead of evaluating on self-contained coding problems, researchers have developed benchmarks focus on solving real-world software engineering issues by operating on an entire coding repository Jimenez et al. [2024a]; Zhang et al. [2023a]; Liu et al. [2023b]. One such popular benchmark is SWE-bench Jimenez et al. [2024a], containing problems where the goal is to modify the repository and resolve a real-world GitHub issue. The authors of SWE-bench have since published a smaller filtered subset of SWE-bench Lite swe [2024], containing 300 total problems, focused on bug fixing issues that only modify a single file in the ground truth patch. In this work, we conduct a detailed classification and analysis of the problems in SWE-bench Lite. We found that some problems lack sufficient information in the problem description to correctly solve the problem. Furthermore, there are also problems containing misleading patches, which can confuse the model. Recognizing these limitations, we further filter SWE-bench Lite to remove such problems and construct SWE-bench Lite-$S$ that can serve as a more rigorous set of problems to evaluate different tools.
Agent-based software development. With the emergence and popularity of agent-based frameworks Xi et al. [2023], recently researchers and industry practitioners have begun developing agent-based approaches to solve software engineering tasks. Devin dev [2024] (and OpenDevin ope [2024b], open-source alternative), is one of the first end-to-end LLM agent-based framework. Devin uses agents to first perform planning based on user requirement, then allows the agent to use file editor, terminal, and web search engine tools to iteratively perform the task. SWE-agent Yang et al. [2024] designs a custom agent-computer interface (ACI) that allows the LLM agent to interact with the repository environment with actions such as reading, editing file, and running bash commands. AutoCodeRover Zhang et al. [2024] is another agent-based approach that provide the LLM agent with specific APIs (e.g., searching methods in certain class) to effectively find the locations that need to be modified to solve the issue. In addition to these highlighted examples, there has been a plethora of other agent-based approaches developed in both open-source Gauthier [2024] and close-source/commercial products Bouzenia et al. [2024]; Chen et al. [2024]; Ma et al. [2024]; lin [2024]; fac [2024]; ibm [2024]; ope [2024a]; mar [2024]; ama [2024]. Compared to these agent-based techniques, Agentless offers a simple and cost-effective solution to tackle real-world software engineering issues. Agentless demonstrates for the first time that an agentless approach can achieve similar performance, without the additional baggage of having to providing excessive tools or modeling complex environment behavior/feedback.

### v2

@@ 5.2 Ablation study on components of Agentless
Next, we examine the impact of using different numbers of sampled candidate patches on the performance of Agentless. Figure 6 shows the number of issues fixed as we increase the number of samples. Note that the sample interval increases by 4 since we use 4 different sets of locations as input. First, we see that by just using 1 greedy sample for each location set, Agentless can already achieve a significant number of correct fixes of 80. We can continue to improve repair performance by adding more samples. However, we observe that the performance plateaus at around 40 samples where adding additional candidate patches does not improve performance. This is because we perform majority voting after test filtering to select the final submission patch, which means that later samples may be ignored since it is difficult for them to offset the majoritively voted patch. Interestingly, we can also see from the figure that if we consider all patch samples (instead of only selecting one patch) for each issue, the total number of possible issues that Agentless can solve is 126 (42.0%). This shows a high upper bound for the potential of Agentless with future work being even better patch re-ranking and selection techniques to further improve the overall performance.
5.2.3 Patch validation ablation.
Table 4: Performance of different patch selection.
Method Performance Avg. \$
Majority voting 77 (25.67%) \$0.00
+Regression test 81 (27.00%) \$0.01
+Reproduction test 96 (32.00%) \$0.25
Finally, we examine the impact of our different test generation and patch selection configurations has on performance. Table 4 shows the result and additional cost of different approaches. We see that by only using majority voting, we can already achieve 77 correct fixes. By adding the existing regression tests, and filter for candidate patches with the lowest amount of regression errors, we can improve performance to 81 issues resolved. Furthermore, the most significant performance improvement was achieved by incorporating additional filtering based on the generated reproduction tests, resulting in the final Agentless performance of 96 fixes. This demonstrates the impact of our patch selection approach, specifically our reproduction test generation, which is able to make use of the high number of candidate patches generated and filter for the correct patch for final submission. However, using reproduction tests also comes with additional costs as Agentless needs to generate these tests which are not provided in the original project repository.
Figure 7: Reproduction test and repair results as number of samples increases.
Next, we look at our reproduction test generation strategy. Figure 7 shows the repair performance (bar, left axis), the total number of selected and plausible reproduction tests (lines, right axis) as we increase the number of candidate reproduction tests generated per issue. Recall that we only select the reproduction test if it can successfully output the Issue reproduced message when evaluated on the original repository. We further consider a selected reproduction test as plausible if it can successfully verify that the ground truth developer patch has fixed the original issue (i.e., output Issue resolved). To begin with, when we only generate one reproduction test (i.e., using the greedy output) per issue, we only produce 100 selected tests for all 300 issues in SWE-bench Lite with the repair performance being 90. We can increase the number of selected reproduction tests by increasing the number of candidate reproduction tests we generate. However, it is interesting to note that the number of plausible reproduction tests does not increase drastically (apart from going from 1 to 5 samples). The reason again is due to some ambiguity in the issue description where it may not contain sufficient information to reproduce and verify the issue has been resolved. Nevertheless, we observe a small improvement in the number of issues resolved, reaching final performance of 96 correct fixes as we increase the number of candidate reproduction tests to 40 per issue.
@@ 6.1 Problem Classification
1) Description quality. We first inspect whether each issue description contains sufficient information to perform the desired task. Figure 8(a) shows the distribution of each category: (i) contains enough information in natural language, (ii) contains reproducible failure example, (iii) contains partially reproducible example, and (iv) does not contain enough information.We observe that while a majority of the tasks in SWE-bench Lite contains sufficient information, with many having some small failure examples to showcase the bug, there is a non-trivial percentage (10.0%) of problems which do not contain enough information. Such problems include those that require implementing a new function with a specific name or adding an error message with a specific string that was not provided in the problem description. [^3] These types of problems still exist in the benchmark despite claims that they have been completely removed by the filtering process according to SWE-bench Lite. This means the test will fail if the function name or error message string does not match exactly, even if the underlying functionality is correctly implemented. Another example of insufficient information are problems that have multiple different interpretations on how to solve the issue, and only a subset of them can pass the ground truth test. For instance, the issue description will outline two possible solutions suggestions with only one of them aligned well with developer intention. Implementing the other proposed solution suggestion will lead to test failure. This highlights the necessity to further sanitize/improve SWE-bench Lite where these problems with uninformative descriptions shall be further excluded.
@@
2) Solution in description. We also examine whether the solution or steps to solve the problem are already provided in the issue description. Figure 8(b) shows the breakdown of our categories: (i) no solution or steps provided, (ii) partial solution provided (e.g., some steps in natural language), (iii) complete solution provided (e.g., complete steps in natural language), (iv) exact patch provided, and (v) misleading solution or steps. Interestingly, we observe that 4.3% of issues contain the exact ground truth patch in the issue description, while an additional 9.7% of issues describe the exact steps required to come up with the correct solution. This shows that certain problems in SWE-bench Lite can be much easier to solve since they provide the solution either in exact code snippets or natural language. Furthermore, we also observe 5.0% of issues contain proposed solution or steps in the issue description that do not reflect the ground truth patch introduced by the developers. This further highlights potential issues with the benchmark, as these discrepancies can mislead tools to generate incorrect solutions simply by following the issue description.
@@
3) Location information. We further check if the issues description contains the correct location information. We divide the granularity into line, function, and file level. Our categories are: (i) exact locations in natural language, (ii) exact locations provided in failure stack traces, iii) related keywords in the issue description that can be used to search for the location, and (iv) not provided. We first observe that only in very few cases (<<10%), the issue provides the exact lines needed to fix the bug. However, this number increases as we increase the granularity to functions and files where we found that around half of the issues already provide the location of the file needed to be edited in the description. To repair a bug or introduce a new feature, finding the location to make the edit is extremely important. As such, we leverage this classification and focus our later analysis on the effect the provided location has on the repair performance of Agentless and baseline approaches.
@@ 6.2 SWE-bench Lite-$S$
CodeStory Aide cod [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o+ [Uncaptioned image] Claude 3.5 S 129 (43.00%) 1 114 (45.78%) 1
Bytedance MarsCode Liu et al. [2024b] [Uncaptioned image] NA 118 (39.33%) 2 106 (42.57%) 2
Honeycomb hon [2024] [Uncaptioned image] NA 115 (38.33%) 3 98 (39.36%) 3
MentatBot men [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o 114 (38.00%) 4 96 (38.55%) 4
Gru gru [2024] [Uncaptioned image] NA 107 (35.67%) 5 94 (37.75%) 5
Isoform iso [2024] [Uncaptioned image] NA 105 (35.00%) 6 91 (36.55%) 6
SuperCoder2.0 sup [2024] [Uncaptioned image] NA 102 (34.00%) 7 87 (34.94%) 7
Alibaba Lingma Agent lin [2024] [Uncaptioned image]
[Uncaptioned image] GPT-4o+
Claude 3.5 S
99 (33.00%) 8 86 (34.54%) 8
Factory Code Droid fac [2024] [Uncaptioned image] NA 94 (31.33%) 10 82 (32.93%) 10
Amazon Q Developer-v2 ama [2024] [Uncaptioned image] NA 89 (29.67%) 12* 76 (30.52%) 13
CodeR Chen et al. [2024] [Uncaptioned image] [Uncaptioned image] GPT-4 85 (28.33%) 14 71 (28.51%) 14*
MASAI Arora et al. [2024] [Uncaptioned image] NA 84 (28.00%) 15 70 (28.11%) 16
SIMA sim [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o 83 (27.67%) 16 71 (28.51%) 14*
IBM Research Agent-101 ibm [2024] [Uncaptioned image] NA 80 (26.67%) 17* 66 (26.51%) 18*
OpenCSG StarShip ope [2024a] [Uncaptioned image] [Uncaptioned image] GPT-4 71 (23.67%) 22 56 (22.49%) 23*
Amazon Q Developer ama [2024] [Uncaptioned image] NA 61 (20.33%) 26 51 (20.48%) 25*
RepoUnderstander Ma et al. [2024] [Uncaptioned image] [Uncaptioned image] GPT-4 64 (21.33%) 25 51 (20.48%) 25*
AutoCodeRover-v2 aut [2024] [Uncaptioned image] GPT-4o 92 (30.67%) 11 79 (31.73%) 11
RepoGraph rep [2024] [Uncaptioned image] GPT-4o 89 (29.67%) 12* 77 (30.92%) 12
Moatless moa [2024]
Claude 3.5 S
80 (26.67%) 17* 67 (26.91%) 17
[Uncaptioned image] GPT-4o 74 (24.67%) 21 62 (24.90%) 21
OpenDevin+CodeAct v1.8 ope [2024b]
Claude 3.5 S
80 (26.67%) 17* 65 (26.10%) 20
Aider Gauthier [2024]
[Uncaptioned image] GPT-4o+
Claude 3.5 S
79 (26.33%) 20 66 (26.51%) 18*
SWE-agent Yang et al. [2024a]
Claude 3.5 S
69 (23.00%) 23 58 (23.29%) 22
[Uncaptioned image] GPT-4o 55 (18.33%) 28 45 (18.07%) 27*
[Uncaptioned image] GPT-4 54 (18.00%) 29 42 (16.87%) 29
AppMap Navie app [2024] [Uncaptioned image] GPT-4o 65 (21.67%) 24 56 (22.49%) 23*
AutoCodeRover Zhang et al. [2024c] [Uncaptioned image] GPT-4 57 (19.00%) 27 45 (18.07%) 27*
RAG Yang et al. [2024a]
Claude 3 Opus
13 (4.33%) 30 10 (4.02%) 30
[Uncaptioned image] GPT-4 8 (2.67%) 32 5 (2.01%) 32
@@
9 (3.00%) 31 6 (2.41%) 31
[Uncaptioned image] GPT-3.5 1 (0.33%) 33 0 (0.00%) 33
Agentless [Uncaptioned image] GPT-4o 96 (32.00%) 9 84 (33.73%) 9
Building on the above problem classifications, we will more rigorously compare and contrast Agentless and existing work. Specifically, we focus on a subset of the problems in SWE-bench Lite after removing the problems that contain the exact patch in the problem description, misleading solutions, or do not provide enough information in the original issue description. This eliminates the less reasonable problems and normalizes the difficulty level of the benchmark. We refer to our subset of 249 problems as SWE-bench Lite-$S$. We note here that our approach of identifying and excluding problematic problems has already been confirmed by a later work done by OpenAI, where they have released a similar filtered benchmark of SWE-bench Verified Chowdhury et al. [2024].
@@
Using the classification results, we further examine the types of problems that are solved by Agentless and prior approaches on SWE-bench Lite-$S$. Figure 9 shows the solve rate of various top-performing open-source and closed-source approaches across the different categories of problems. We first examine if having code examples to reproduce the error in the issue description can help the LLM better solve the issue in Figure 9(a). Surprisingly, we found that the solve rate of all prior approaches drop when evaluated on the problems with reproducible code examples. Many agent-based approaches Yang et al. [2024a]; ope [2024b]; Chen et al. [2024] attempt to first reproduce the error, however, this may not improve performance even on problems with already provided reproducible examples. However, we observe the performance for Agentless remains very high on the problems with reproducible code examples. This is because Agentless generates reproduction tests using the original issues descriptions, hence can better make use of the reproducible code examples provided (similarly for MentatBot which also contains an explicit test generation step). This demonstrate the importance of the test generation stage for patch selection. Next, we look at the effect of ground truth patch/solution in the issue description. Figure 9(b) shows the expected result where all selected techniques perform better on issues that already provide solution steps in natural language. Furthermore, in Figure 9(c), we examine the solve rate with respect to the location information provided in the issues description. Unsurprisingly, we found that the highest solve rates are on problems where the location is provided in natural language followed by stack traces. The most difficult problems are those that do not contain any clues about the location of the issue in the description. We observe that compared with closed-source approaches, Agentless performs comparably when the location is provided in either natural language, stack trace, or keywords. However, the closed-source agent tools perform better compared to Agentless in the case where no location clue is provided. This highlights an advantage of agent-based tools in solving these more complex problems where they are able to use complex code search tools. This represents potential future work for Agentless to target and further improve these types of problems.
6.3 SWE-bench Verified
Table 6: Performance on SWE-bench Verified.
Tool LLM % Resolved
Tools too [2024] [Uncaptioned image]
Claude 3.5 Sonnet (20241022)
245 (49.00%)
Claude 3.5 Haiku
203 (40.60%)
Solver sol [2024] [Uncaptioned image] NA 227 (45.40%)
Gru gru [2024] [Uncaptioned image] NA 226 (45.20%)
Honeycomb hon [2024] [Uncaptioned image] NA 203 (40.60%)
Amazon Q Developer-v2 ama [2024] [Uncaptioned image] NA 194 (38.80%)
Factory Code Droid fac [2024] [Uncaptioned image] NA 185 (37.00%)
Amazon Q Developer ama [2024] [Uncaptioned image] NA 128 (25.60%)
Composio SWEkit com [2024]
Claude 3.5 Sonnet
203 (40.60%)
AutoCodeRover-v2 aut [2024] [Uncaptioned image] GPT-4o 192 (38.40%)
SWE-agent Yang et al. [2024a]
Claude 3.5 Sonnet
168 (33.60%)
[Uncaptioned image] GPT-4o 116 (23.20%)
[Uncaptioned image] GPT-4 112 (22.40%)
Alibaba Lingma Agent lin [2024] Lingma SWE-GPT 72b 144 (28.80%)
AppMap Navie app [2024] [Uncaptioned image] GPT-4o 131 (26.20%)
Agentless [Uncaptioned image] GPT-4o 194 (38.80%)
Similar to SWE-bench Lite-$S$ and inspired by similar concerns in Section 6.1, OpenAI has produced a newly filtered dataset of SWE-bench Verified Chowdhury et al. [2024] validated by human developers to ensure each issue has sufficient amounts of information to be solved. Table 6 shows the performance of Agentless compared with prior agent-based approaches on SWE-bench Verified. Similar to the results on SWE-bench Lite, Agentless maintains its strong performance and is able to solve 194 out of 500 problems (38.80%). Agentless is able to achieve the second highest performance compared with open-source approaches and perform better than many closed-source / commercial techniques. Furthermore, we note that Agentless performs the best among all techniques that use GPT-4o as the LLM.
7 Threats to Validity
Internal. One threat to validity comes from the data leakage of ground truth developer patches in SWE-bench Lite being part of the training data for GPT-4o. Since GPT-4o is a closed-source model, we do not have access to the training data. Meanwhile, we note here that prior work almost exclusively used similar closed-source LLMs (e.g., GPT-4o, GPT-4, Claude-3.5, etc), and our approach can outperform all existing open-source solutions with same models. Furthermore, the authors of SWE-bench Jimenez et al. [2024a] compared the resolve rate of issues collected before and after the knowledge cutoff date of GPT-4, and did not find any significant difference. To completely address this threat, we would need to retrain GPT-4o from scratch which would be infeasible for an academic project.
External. One main external threat comes from our evaluation dataset of SWE-bench Lite. While the performance of Agentless might not generalize to other datasets, SWE-bench Lite is by far the most popular evaluation dataset which contains a diverse range of problems. In addition, OpenAI has independently performed an extensive evaluation of Agentless and other open-source solutions on SWE-bench Lite, SWE-bench, and their new SWE-bench Verified benchmark, further confirming that Agentless outperforms all other open-source agents Chowdhury et al. [2024]. Moreover, on Sept. 12th 2024, OpenAI just released the new o1 model family and also adopted Agentless as the top approach to showcase their performance on SWE-bench OpenAI [2024c]. In the future, we plan to further address this threat by evaluating Agentless on other benchmarks.